determine the amount of heat (in kj) given off when 4.53 * 10^4 g of ammonia is produced according to the equation below. assume that the reaction takes place under standard - state conditions at 25 celsius.

N2(g) + 3H2(g) ---> 2NH3

?H= -92.6 kJ/mol

enter the answser in scientific notation
Answer #1

Molar mass of ammonia , NH3 is = At.mass of N + (3xAt.mass of H )

                                                   = 14 + (3x1)

                                                   = 17 g/mol

N2(g) + 3H2(g) ---> 2NH3 ; \Delta H= -92.6 kJ/mol

According to the above balanced equation ,

for the formation of 2 moles of NH3 the heat evolved is 92.6 kJ

                                         OR

for the formation of 2x17 g of NH3 the heat evolved is 92.6 kJ

for the formation of 4.53x104 g of NH3 the heat evolved is Z kJ

Z=\frac{92.6}{2\times 17}\times 4.53\times 10^{4}=1.23\times 10^{5}kJ

Therefore the amount of heat (in kj) given off is 1.23x105 kJ